# Scaling Multi-Agent Epistemic Planning through GNN-Derived Heuristics

### v3

@@ 2.2. Multi-Agent Epistemic Planning
To the best of our knowledge, the most accepted formalization of a comprehensive action language for multi-agent epistemic planning is (baral2021action). Note that other languages capable of reasoning about DEL also exist (muise2015planning; KR2021-12), but they limit their expressiveness in favor of efficiency and are therefore not considered here.
@@ 3.1. Training Data Generation
The next challenge we addressed was designing a suitable data representation for integration within a GNN. Each e-state is formalized as a pointed Kripke structure (Definition 2.1), namely
@@
Each world is assigned a symbolic integer identifier through a mapping Intuitively, this assigns consecutive integers to newly observed worlds. However, since this assignment depends on the order in which worlds are generated, making not invariant across runs, leading to inconsistencies between datasets. This representation has the lowest computational burden for data generation, training, and inference, but it incurs the greatest information loss, as similarities between worlds are not preserved.
@@ 3.2.1. Goal Encoding within the State Embedding
In the resulting formulation, two types of nodes are encoded for each e-state embedding: (i) state nodes, corresponding to worlds in the epistemic state; (ii) goal nodes, corresponding to symbols describing the goal conditions. This is exemplified in Figure 1, where the left portion of the expanded hashed-based e-state (in teal) shows the goal encoding, and the right portion (in magenta) depicts the e-state itself. The two are connected via a shared graph structure using nodes and edges that use constants and special identifiers throughout the process.
@@ 3.4. Training Neural Distance Estimator
Bitwise embedding of node identifiers and edge attributes: Each node bit-vector is projected through a two-layer MLP, producing dense embeddings of dimension . Similarly, each edge attribute is processed through yielding -dimensional embeddings. This bitwise encoding provides a compact and lossless representation of node identifiers while supporting continuous optimization and stable gradients.
@@
Graph-level summarization: The node embeddings from the final GINE layer are aggregated into a fixed-size representation via mean pooling: This operation is permutation-invariant and scale-independent, ensuring stable representations across graphs of different sizes. In contrast, sum pooling scales with , potentially biasing larger graphs, while attention pooling introduces learnable weights but increases variance and computational cost. Mean pooling therefore offers the best generalization–variance trade-off for distance regression.
@@ 4.2. Experimental Results
Finally, Table 5 presents aggregated results comparing our approach with existing heuristics in -EFP. For this evaluation, we use -, as it represents the most general-purpose configuration. A detailed comparison with individual heuristics (C_PG, L_PG, S_PG, and SUB) is provided in Appendix G. We note that the only metric with significant interpretability is the number of instances solved, since the -EFP heuristics solve a highly diverse set of instances, making other aggregated measures less informative.
@@ 4.3. Discussion
GNN demonstrates informative and robust performance across a range of planning domains, as evidenced by the aggregate metrics reported in Table 3. With the exception of the GR domain, GNN consistently reduces the number of explored nodes compared to uninformed search, highlighting the efficiency gains enabled by the learned heuristic.
@@
Table 5 further compares GNN against individual heuristics within -EFP. GNN performs on par with the best individual heuristics, demonstrating its strength and viability as an alternative approach. At the same time, we view GNN-based heuristics as a complementary tool to enhance solver coverage, for example through integration into portfolio-based approaches, rather than as a replacement for existing heuristics.
@@ 5. Limitations and Future Directions
Finally, an important next step is to integrate our GNN-based heuristics estimates into more advanced and potentially more effective search frameworks, such as MCTS sutton1998reinforcement. We believe this direction holds substantial promise for improving the scalability and adaptability of MEP solvers.

### v4

@@ 2.2. Multi-Agent Epistemic Planning
To the best of our knowledge, the most widely accepted formalization of a comprehensive action language for multi-agent epistemic planning is (baral2021action). Note that other languages capable of reasoning about DEL also exist (muise2015planning; KR2021-12), but they limit their expressiveness in favor of efficiency and are therefore not considered here.
@@ 3.1. Training Data Generation
The next challenge we address is designing a suitable data representation for integration within a GNN. Each e-state is formalized as a pointed Kripke structure (Definition 2.1), namely
@@
Each world is assigned a symbolic integer identifier through a mapping Intuitively, this assigns consecutive integers to newly observed worlds. However, since this assignment depends on the order in which worlds are generated, is not invariant across runs, leading to inconsistencies between datasets. This representation has the lowest computational burden for data generation, training, and inference, but it incurs the greatest information loss, as similarities between worlds are not preserved.
@@ 3.2.1. Goal Encoding within the State Embedding
In the resulting formulation, two types of nodes are encoded for each e-state embedding:
state nodes, corresponding to worlds in the epistemic state;
goal nodes, corresponding to symbols describing the goal conditions.
This is exemplified in Figure 1, where the left portion of the expanded hashed-based e-state (in teal) shows the goal encoding, and the right portion (in magenta) depicts the e-state itself. The two are connected via a shared graph structure using nodes and edges that use constants and special identifiers throughout the process.
@@ 3.4. Training Neural Distance Estimator
Bitwise embedding of node identifiers and edge attributes: Each node bit-vector is projected through a two-layer MLP,
producing dense embeddings of dimension . Similarly, each edge attribute is processed through
yielding -dimensional embeddings. This bitwise encoding provides a compact and lossless representation of node identifiers while supporting continuous optimization and stable gradients.
@@
Graph-level summarization: The node embeddings from the final GINE layer are aggregated into a fixed-size representation via mean pooling:
This operation is permutation-invariant and scale-independent, ensuring stable representations across graphs of different sizes. In contrast, sum pooling scales with , potentially biasing larger graphs, while attention pooling introduces learnable weights but increases variance and computational cost. Mean pooling therefore offers the best generalization–variance trade-off for distance regression.
@@ 4.2. Experimental Results
Table 5 compares our approach with existing heuristics in -EFP. For this evaluation, we use -, as it represents the most general-purpose configuration. A detailed comparison with individual heuristics (C_PG, L_PG, S_PG, and SUB) and their description is provided in Appendix G. We note that the only metric with significant interpretability is the number of instances solved, since the -EFP heuristics solve a highly diverse set of instances, making aggregated measures less informative.
@@ 4.3. Discussion
GNN demonstrates informative and robust performance across a range of planning domains, as evidenced by the aggregate metrics reported in Table 3. With the exception of the GR domain, GNN consistently reduces the number of explored nodes compared to uninformed search, highlighting its efficiency gains.
@@
[…]
@@ 5. Limitations and Future Directions
Finally, an important next step is to integrate our GNN-based heuristic estimates into more advanced search frameworks, such as Monte Carlo Tree Search sutton1998reinforcement. We believe this approach has strong potential to improve the scalability and adaptability of MEP solvers.